TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes


Created by Main.UnknownUser on 2006-11-30 - 20:07

preparation for cool down sec 78 - Nov. 30

Microsoft Excel Spreadsheet Sector_78_Planning_60A.xls:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x30x20x01/Sector_78_Planning_60A.xls

PDF RAT_20061130.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x30x20x01/RAT_20061130.pdf

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 30.11.06 Non conformity: 6 temperature sensor cables need repair (A. Suraci)
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX) -- boris.bellesia - 30 Nov 2006 - 20:05

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-29 - 18:50

Preparation for Cool Down sec.78 - Nov 29

PDF RAT_20061129.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x29x18x25/RAT_20061129.pdf

Mpp Sector_78_Preparation_of_Cool_Down_w48.mpp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x29x18x25/Sector_78_Preparation_of_Cool_Down_w48.mpp
UPDATED PLANNING (comments are welcome!)

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 3 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 29.11.06 Definition of owner of activity and pumping needs for current leads insulation vacuum (DFBX)
  • 29.11.06 Pumping of insulation vacuum for DFBX
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-29 - 12:57

RAT at Point 5, Nov 29

PDF RAT_20061129.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x29x12x54/RAT_20061129.pdf

Comments:

Open Issues

  • 29.11 WIFI problems
  • 29.11 ELETTAS threshold setup
  • 23.11 Ripples in RQT12R5B1 power converter.
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-28 - 16:40

Preparation for Cool Down sec. 78 - Nov 28

PDF RAT_20061128.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x28x16x37/RAT_20061128.pdf

-- Main.MircoCoccoli - 28 Nov 2006 - 16:38

Open Issues:

AC non-conformity

  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair-Wed.29th)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG, DFBMH
  • 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
  • 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC, DFBAN (Flohe)
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (A.Suraci: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: (done: DFBAO; next(w48): DFBAN, DFBMA, DFBMC, DFBMH, DFBX))
  • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done; next: DFBMA, DFBMC, DFBX)
Safety General remark: written communication in advance to announce operations!!!
  • 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-27 - 19:08

RAT - Preparation for Cool Down sec.78 - Nov 27

PDF RAT_20061127.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x27x19x02/RAT_20061127.pdf

Main.MircoCoccoli - 27 Nov 2006 - 19:02

Open Issues:

AC non-conformity

  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83 (P.Chambouvet: about 2 weeks of work starting Mon 27 Nov)
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 25.11.06 Functional test on the leads heaters: DFBMA, DFBMC, DFBX
  • 25.11.06 120A and 600A positioning on DFBAN and DFBMH
  • 23.11.06 Cable temporary positioning on the transport side of the tunnel (A.Vergara: next year)
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 20.11.06 Galvanic insulation installation on all DFBs
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected (D.Bozzini: 80% done)
  • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
DFB Commissioning
  • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: ongoing according to schedule (next: DFBMA, DFBMC, DFBX); in week 48 will begin ElQA-TP4-B, i.e. HV tests)
  • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done; next: DFBMA, DFBMC, DFBX)
Safety General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J.C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 23.11.06 Planning of consignations/deconsignations
  • 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-27 - 11:34

RAT at Point 5, Nov 27

PDF RAT_20061127.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x27x11x29/RAT_20061127.pdf

Comments:

Open issues

  • 23.11 Ripples in RQT12R5B1 power converter.
  • 23.11 Cables inversion on rack RYMCA04.
  • 20.11 Problem with water conductivity in DQS: reason known, insulation test still to be performed.
  • 13.11 DC cables high voltage test (RR57 done. Waiting UJ56)
  • 13.11 Test de circulation ED (waiting UJ56)
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-11-24 - 20:37

Preparation Cool-Down Sector 78-RAT 24th Nov

PDF RAT_20061124.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x24x20x25/RAT_20061124.pdf

Comments: Open Issues:

  • AC non-conformity
    • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83 (P.Chambouvet: about 2 weeks of work starting Mon 27 Nov)
    • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads) (A.Suraci: 4 cables are in repair)
    • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
    • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
  • DC cabling
    • 23.11.06 Cable temporary positioning on the transport side of the tunnel (A.Vergara: next year)
    • 22.11.06 Pre-connection of 120A cables in LSSR7
    • 20.11.06 Galvanic insulation installation on all DFBs
    • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
    • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • Assembly -
  • Vacuum -
  • Cryogenics
    • 22.11.06 Cryo-valves remote control in CCC by AB/CO
    • 22.11.06 Control of quench valve
    • 20.11.06 DFB instrumentation cables to be connected (D.Bozzini: 80% done)
    • 17.11.06 Instrumentation cable HV tests in LSS8L (A.Suraci: 4 cables to be tested)
  • DFB Commissioning
    • 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: ongoing according to schedule; next week will begin ElQA-TP4-B, i.e. HV tests)
    • 17.11.06 Pumping of insulation vacuum for all DFBs between 7 and 8 (P.Chambouvet: DFBAN done)
  • Safety
    • General remark: written communication in advance to announce operations!!!
    • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
    • 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J.C.Guillaume and approval by J.Etheridge.
    • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
    • 23.11.06 Planning of consignations/deconsignations
    • 16.11.06 Cool down safety procedure & access conditions
  • Preparation of powering test
    • 22.11.06 Pretest of 60A converters in sector 78, location C16L8. (B.Bellesia: stopped due to LQA hardware problems; expected end by Tuesday 28 Nov)
    • 22.11.06 Installation of flexible cables by CV (UA83)
    • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
  • AOB
    • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
    • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.BlancaPereaSolano - 24 Nov 2006 - 20:29

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-23 - 15:23

Preparation for Cool-Down Sec78 - Thu Nov, 23rd

PDF RAT_20061123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x23x14x53/RAT_20061123.pdf

Comments:

    • AC non-conformity
  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.

    • DC cabling
  • 23.11.06 Cable temporary positioning on the transport side of the tunnel
  • 23.11.06 Adjustment of cable length on DFBAO
  • 22.11.06 Missing labels on DFBAO instrumentation cables
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 20.11.06 Galvanic insulation installation of 600 A cables to the DFBAO
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
    • Assembly: -

    • Vacuum: -

    • Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected
  • 17.11.06 Instrumentation cable HV tests in LSS8L

    • DFB Commissioning
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum DFBAN, DFBAO, DFBMC, DFBMA, DFBXG

    • Safety: -

    • General remark: written communication in advance to announce operations!!!
  • 23.11.06 Wooden structure for UA access restriction (that will allow transport).
  • 23.11.06 Preparation of a prototype plastic protection or the monorail (protection derived from a plastic tube) by J.C.Guillaume and approval by J.Etheridge.
  • 23.11.06 Water leak on the tunnel concrete wall to be fixed (33L8).
  • 23.11.06 Planning of consignations/deconsignations
  • 16.11.06 Cool down safety procedure & access conditions

    • Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.

    • AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

-- Main.MircoCoccoli - 23 Nov 2006 - 14:55

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-11-23 - 12:20

Preparation for Cool-Down RAT - Tuesday, November 21st

PDF RAT_20061121.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x21x12x15/RAT_20061121.pdf

Open Issues

  • AC non-conformity
    • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83
    • 21.11.06 Non-conformity on instrumentation cables (LT, PT, Level Gauges, …)
    • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
    • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
    • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • DC cabling
    • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
    • 17.11.06 600 A cable positioning on DFBA
    • 20.11.06 Connections (made by TS/MME) of 120 A cables to be verified (TS/EL)
    • 20.11.06 Connections of 600 A cables to the DFBAO to be verified
  • Assembly

  • Vacuum
    • 20.11.06 21 leak detectors to be placed all along the arc 78
  • Cryogenics
    • 20.11.06 DFB instrumentation cables to be connected
    • 20.11.06 Leaks on DFBAO gas recovery line in repair/installation of helicoflex
    • 17.11.06 Instrumentation cable HV tests in LSS8L
  • DFB Commissioning
    • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
    • 13.11.06 ElQA -TP4-A in the DFBAO, DFBAN, DFBAX
    • 17.11.06 Pumping of insulation vacuum DFBAN, DFBAO, DFBMC, DFBMA, DFBXG
  • Safety
  • General remark: written communication in advance to announce operations!!!
    • 16.11.06 Cool down safety procedure & access conditions
    • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
  • Preparation of powering test
    • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be cleaned!
  • AOB
    • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
    • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-23 - 11:53

RAT at Point5, Nov 23

PDF RAT_20061123.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x23x11x08/RAT_20061123.pdf

Comments:

Next RAT on Monday 27

Open Issues

  • 20.11 Problem with water conductivity in DQS
  • 13.11 DC cables high voltage test (RR57 done. Waiting UJ56)
  • 13.11 Ventilation in UJ56.
  • 13.11 Test de circulation ED (waiting UJ56)
  • 20.11 Water conductivity in DQS – checked tomorrow (23.11)
  • 22.11 DQS acousting shielding and cabling – No news
  • 22.11 Full power ventilation test in RR57 – Friday 8th Dec. 8:30-12:00
  • 22.11 Fire detection – awaiting certification
  • 22.11 Wi-Fi damaged cable has to be repaired – no news

Edit, attach file or add comment


Created by Main.MircoCoccoli on 2006-11-22 - 16:19

Preparation for Cool-Down of Sector 78, 22 Nov

PDF RAT_20061122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x22x16x14/RAT_20061122.pdf

Comments:

AC non-conformity

  • 21.11.06 Transport and installation of the heater system of the current leads for RR77 and UA83
  • 21.11.06 Non-conformity on instrumentation cables (temperature sensors on current leads)
  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • 20.11.06 Galvanic insulation installation of 600 A cables to the DFBAO
  • 22.11.06 Pre-connection of 120A cables in LSSR7
  • 22.11.06 Missing labels on DFBAO instrumentation cables
Assembly - Vacuum - Cryogenics
  • 22.11.06 Cryo-valves remote control in CCC by AB/CO
  • 22.11.06 Control of quench valve
  • 20.11.06 DFB instrumentation cables to be connected
  • 17.11.06 Instrumentation cable HV tests in LSS8L
DFB Commissioning
  • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum DFBAN, DFBAO, DFBMC, DFBMA, DFBXG
Safety General remark: written communication in advance to announce operations!!!
  • 16.11.06 Cool down safety procedure & access conditions
  • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
Preparation of powering test
  • 22.11.06 Pretest of 60A converters in sector 78, location C16L8.
  • 22.11.06 Installation of flexible cables by CV (UA83)
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-22 - 15:45

Last RAT meeting at Point5 - Wednesday 22

PDF RAT_20061122.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x22x15x44/RAT_20061122.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-11-20 - 19:55

Preparation cool-down of Sector 78, 20 Nov

PDF Sector_78_Preparation_of_Cool_Down_w47.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x20x19x42/Sector_78_Preparation_of_Cool_Down_w47.pdf

PDF RAT_20061120.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x20x19x42/RAT_20061120.pdf

Mpp Sector_78_Preparation_of_Cool_Down_w47.mpp:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x20x19x42/Sector_78_Preparation_of_Cool_Down_w47.mpp

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converetrs & VAC.
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
DC cabling
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • 13.11.06 120 A cable connections in DFBAO, DFBMC, DFBMA and DFBX LSS8L
  • 17.11.06 600 A cable positioning on DFBA
  • 20.11.06 Connections (made by TS/MME) of 120 A cables to be verified (TS/EL)
  • 20.11.06 Connections of 600 A cables to the DFBAO to be verified
  • 20.11.06 Flohe on DFBAO or DFBAM? To be decided

Assembly - Vacuum

  • 20.11.06 21 leak detectors to be placed all along the arc 78
  • 17.11.06 Bellows damaged in Q6-Q7 LSS8L
Cryogenics
  • 20.11.06 DFB instrumentation cables to be connected
  • 20.11.06 leaks on DFBAO in reparation
  • 17.11.06 Instrumentation cable HV tests in LSS8L
  • 17.11.06 Pressurized air supply to valve on WRL
  • 13.11.06 Support on DFBAO WRL
DFB Commissioning
  • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum will start next week in 7R
Safety General remark: written communication in advance to announce operations!!!
  • 16.11.06 Cool down safety procedure & access conditions
  • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
  • 17.11.06 Monorail in Point 8 locked off on Monday (20.11) starting from 10.30!!!
Preparation of powering test
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be cleaned!
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX) -- Main.bellesia - 20 Nov 2006 - 19:53

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-20 - 16:21

Last RAT meeting in pint 5, Monday Nov 20.

Please find enclose the minutes of our last RAT meeting in point 5.

Regards

PDF RAT_20061120.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x20x16x18/RAT_20061120.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-11-20 - 15:35

Warm Magnets - Satefy

A well protected magnet. And a non protected magnet.

JPEG Image044.jpg:
Image044.jpg
A non so well protected magnet
JPEG Image042.jpg:
Image042.jpg
A well protected magnet

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-11-20 - 10:33

tests en eau des cables refroidis RR57 du 20 novembre: TOUT est OK

JPEG tests_en_eau_cables_refoid_RR57_.jpg:
tests_en_eau_cables_refoid_RR57_.jpg

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-17 - 16:54

Preparation of cool-down in sector 78, 17 Nov

PDF RAT_20061117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x17x16x48/RAT_20061117.pdf

Comments:

AC non-conformity

  • 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current)
  • 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converetrs & VAC.
  • 13.11.06 Non-conformity of the DC cabling of the orbit corrector power converters
DC cabling
  • 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA and DFBMC
  • 13.11.06 120 A cable connections in DFBAO, DFBMC, DFBMA and DFBX LSS8L
  • 17.11.06 600 A cable positioning on DFBA
Assembly -

Vacuum

  • 17.11.06 Leak in arc (Q11-8L)
  • 17.11.06 Bellows damaged in Q6-Q7 LSS8L
Cryogenics
  • 17.11.06 Instrumentation cable HV tests in LSS8L
  • 17.11.06 Cryogenic valve manipulation on DFBX
  • 17.11.06 Pressurized air supply to valve on WRL
  • 13.11.06 Support on DFBAO WRL
DFB Commissioning
  • 17.11.06 Prototype of galvanic insulation to be tested at Point 6
  • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
  • 17.11.06 Pumping of insulation vacuum will start next week in 7R
Safety General remark: written communication in advance to announce operations!!!
  • 16.11.06 Cool down safety procedure & access conditions
  • 16.11.06 Definition of Safety procedure for Powering Procedure Test (aka Dry Runs)
  • 17.11.06 Monorail in Point 8 locked off on Monday starting from 10.30!!!
Preparation of powering test
  • 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be cleaned!
AOB
  • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
  • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-11-17 - 14:41

Essais en court-circuits UJ56 et RR57

Bonjour,

Les visites des installations électriques des ouvrages cités en objet ont été réalisées à ce jour. Aucune non conformité ne s'oppose à ce que les essais en court-circuit soient effectués à partir du 20 novembre 2006. Les rapports de visite parviendront ultérieurement aux responsables de chaque installation.

Cordialement,

Caroline BENOIT-GONIN Norisko Equipements pour SC/GS Office 57-2-024 caroline.benoit-gonin@cernNOSPAMPLEASE.ch

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-17 - 12:09

Test on the DQSQD switch, today Friday Nov 17.

Please find enclose a small report of the tests held in UA63 concerning the 13kA switch DQSQD.

Regards

PDF RAT_20061117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x17x12x07/RAT_20061117.pdf

Comments:

Edit, attach file or add comment


Created by Main.MirkoPojer on 2006-11-16 - 17:10

Preparation of Cool-Down in Sector 78, 16 Nov

PDF RAT_20061116.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x16x17x00/RAT_20061116.pdf

Comments:

  • Open Issues
    • 13.11.06 Cryo. Valves manipulation on the DFBMH, DFBAN and DFBAO
    • 13.11.06 Non-conformity of the mini-crates of cryo instrumentation
    • 13.11.06 Cabling non-conformity affecting QPS and 60A Orbit Correctors IST
    • 13.11.06 WRL DFBAO missing support LSS8L
    • 13.11.06 3 x 600 A power cable positioning LSS8L
    • 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7
    • 13.11.06 ElQA-TP4-A in the DFBAO, DFBAN, DFBAX
    • 13.11.06 LSS8L Q6-Q7 bake-out
    • 13.11.06 120 A Cables connection in DFBMC, DFBMA and DFBX LSS8L
    • 13.11.06 6 kA and 13 kA cables positioning at DFBMA and DFBMC
    • 13.11.06 Quench protection continuity tests and cables assignment LLS7R (DFBMH, DFBAN) and LSS8L (DFBMC, DFBMA and DFBX)
    • 13.11.06 DS7R and DS8L leak test
    • 13.11.06 Leak in R7
    • 13.11.06 Galvanic Insulation 600 A, 6 kA, 13 kA for LSS8L and LSS7R available
    • 15.11.06 Leak in arc
    • 15.11.06 Leak test to be performed after the replacement of bellows

-- Main.MirkoPojer - 16 Nov 2006 - 17:52

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-16 - 14:27

Event analysis, DQSQD switch behavior during the AUG simulation on Friday Nov 3

Please find enclose a short description of the events related to the behavior of the DQSQD 13kA switch during the AUG simulation on Friday Nov 3.

Regards Rosario

PDF MinutesFailureEE_20061117.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x16x14x16/MinutesFailureEE_20061117.pdf

Comments:

Edit, attach file or add comment


Created by Main.UnknownUser on 2006-11-15 - 16:05

Preparation cool-down of Sector 78, 15 Nov

PDF RAT_20061115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x15x14x56/RAT_20061115.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-15 - 12:00

RAT today in point 5, Nov 15

Please find enclose the last RAT minutes. Meeting held today, Wed Nov 15, in point 5.

Main issue: schedule in UJ56 and RR57.

Regards

PDF RAT_20061115.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x15x11x52/RAT_20061115.pdf

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-11-14 - 17:45

Preparation of the Cool-Down in Sector 78, 14 Nov.

PDF RAT_20061114.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x14x17x42/RAT_20061114.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-11-14 - 16:37

RR77, ca avance...

lundi 13 , tests des cables refroidis sans eau?...tout est OK....

JPEG cable_refroidi_RR57JPG.JPG:
cable_refroidi_RR57JPG.JPG

Comments:

Edit, attach file or add comment


Created by Main.BlancaPereaSolano on 2006-11-13 - 17:55

Preparation Cool-Down Sector 78-RAT 13th Nov

PDF RAT_20061113.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x11x13x17x54/RAT_20061113.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-13 - 16:35

RAT meeting, Mon Nov 13

Please find enclose the minutes of the last RAT meeeting held this morning, Nov 13, in point 5.

Main issue: the readiness of the RR57 and UJ56, DC cables, elettas, etc..

Regards

PDF RAT_20061113.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x13x16x33/RAT_20061113.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-08 - 16:10

RAT in point 5, Nov 8

Please find enclose the minutes of the last RAT meeting, held this morning in point 5.

Main issues:

- Preparation of the SCT in RR57 and UJ56

- Last tests in UA63

Reg

PDF RAT_20061108.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x11x08x16x08/RAT_20061108.pdf

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-11-06 - 15:25

The Polar Express

Polarity guys in action

JPEG Image021.jpg:
Image021.jpg

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-03 - 16:18

Last RAT this morning, Nov 3, in point 6

Please find enclose the RAT meeting minutes of our last meeting in point 6.

Main issues: a) results of the 24-h heat run b) UPS test c) AUG simulation

Regards

PDF RAT_20061103.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x03x16x16/RAT_20061103.pdf

Comments:

Edit, attach file or add comment


Created by Main.JeanPierreMalodDognin on 2006-11-03 - 15:39

RR57 en finition

JPEG Tests_cables_refroidis_RR57.jpg:
Tests_cables_refroidis_RR57.jpg

JPEG IMGP8910.JPG:
IMGP8910.JPG

JPEG Redimensionnement_de_IMGP8903.JPG:
Redimensionnement_de_IMGP8903.JPG

JPEG IMGP8938.JPG:
IMGP8938.JPG

JPEG IMGP8950.JPG:
IMGP8950.JPG

JPEG IMGP8944.JPG:
IMGP8944.JPG

Comments:

Edit, attach file or add comment


Created by Main.AntonioVergara on 2006-11-02 - 18:02

RAT meeting 02-11-2006

PDF RAT_061102.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x02x18x01/RAT_061102.pdf

Comments:

Edit, attach file or add comment


Created by Main.RosarioPrincipe on 2006-11-01 - 10:11

RAT meeting Nov 1st in CCC

Please find enclose the minutes of the last RAT meeting held today in the CCC.

Remember 24-h heat run today in UA6.

Regards.

PDF RAT_20061101.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x01x10x09/RAT_20061101.pdf

Comments:

Edit, attach file or add comment


Created by Main.JacekSzkutnik on 2006-11-01 - 09:22

RAT meeting this morning in point 6

Please find enclose the minutes of our last RAT meeting, held this morning in SX6.

Main issue: the 24-h heat run scheduled from Wed 01.11 to Thu 02.11.

Regards

PDF RAT_20061031.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint6x2006x11x1x09x21/RAT_20061031.pdf

Comments:

Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2006-11-01 - JacekSzkutnik
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login